\begin{thebibliography}{99}

\bibitem{CKW} A.~Conca, C.~Krattenthaler, and J.~Watanabe. 
\emph{Regular sequences of symmetric polynomials.} Rend. Semin. Math. Univ. 
Padova {\bf 121}, 179--199 (2009). 

\bibitem{CoCoA}
The CoCoATeam. \emph{CoCoA: a system for doing Computations in
Commutative Algebra.} Available at \texttt{http://cocoa.dima.unige.it.}

\bibitem{DZ-2} R.~Dvornicich and U.~Zannier. \emph{Newton functions generating symmetric 
fields and irreducibility of {S}chur polynomials.} Adv. Math. {\bf 222}, no. 6, 1982--2003 (2009).


\bibitem{DE} D.~Eisenbud. \emph{Commutative Algebra.} With a view toward algebraic 
geometry. Graduate Texts in Mathematics, {\bf 150} Springer-Verlag, New York, xvi+785 pp. (1995).

\bibitem{Froberg-Shapiro} R.~Fr\"oberg and B.~Shapiro. \emph{Vandermonde varities and 
relations among Schur polynomials.} arXiv:1302.1298v1 [math.AC].

\bibitem{Ha-Wa} T.~Harima and J.~Watababe. \emph{The strong {L}efschetz property for 
{A}rtinian algebras with non-standard grading.} J. Algebra. { \bf 311}, no. 2, 511--537 (2007).

\bibitem{Hos} M.~Hochster. \emph{Big Cohen-Macaulay modules and algebras and embeddability 
in rings of Witt vectors.} Queen’s Papers on Pure and Applied Math. { \bf 42}, 106--195 (1975).

\bibitem{KM} N.~Kumar and I.~Martino. \emph{Regular sequences of power sums and 
complete symmetric polynomials.} Matematiche (Catania). {\bf 67}, no. 1, 103--117 (2012).

\bibitem{Lam} T. Y.~Lam and K.H.~Leung. \emph{On Vanishing sums of roots of unity.} 
Journal of algebra. {\bf 224 }, no. 1, 91--109  (2000).

\bibitem{I-G} I. G.~Macdonald. \emph{Symmetric functions and Hall polynomials.} 
Second edition. With contributions by A. Zelevinsky. Oxford Mathematical Monographs. 
Oxford Science Publications. The Clarendon Press, Oxford University Press, 
New York, x+475 pp, (1995). 

\bibitem{Mann} H. B.~Mann. \emph{On linear relations between the roots of unity.}
Mathematika. {\bf 12}, 107--117 (1965).


\bibitem{Milne} J. S.~Milne. \emph{Algebraic geometry.} Allied Publishers. (2005).


\bibitem{Poonen-Rub} B.~Poonen and M.~Rubinstein. \emph{The number of intersection points 
made by the diagonal of a regular polygon.} SIAM J. Discrete Math. { \bf 11}, no. 1, 135--156 (1998).


\bibitem{RS} R. P.~Stanley. \emph{Weyl groups, the hard {L}efschetz theorem, and the {S}perner
              property.} SIAM J. Algebraic Discrete Methods. {\bf 1}, no. 2, 168--184 (1980).

\bibitem{AW} A.~Wiebe. \emph{The {L}efschetz property for componentwise linear ideals and
              {G}otzmann ideals.} Comm. Algebra. {\bf 32}, no. 12, 4601--4611 (2004).
              
\end{thebibliography}

\documentclass[reqno, a4paper, 11pt]{amsart}
\usepackage[utf8x]{inputenc}
\usepackage[dvips]{epsfig}
\usepackage{amsgen, amstext,amsbsy,amsopn,amssymb, amsthm, 
mathptmx, amsfonts,amssymb,amscd,amsmath,nicefrac,euscript,enumerate,url,verbatim,calc,framed}
\usepackage[all]{xy}

\DeclareMathOperator{\dett}{det}

\DeclareMathOperator{\tensor}{\otimes}

\DeclareMathOperator{\pnt}{\raise 0.5mm \hbox{\large\bf.}}

\newtheorem{thm}{\bf Theorem}[section]
\newtheorem{lem}[thm]{\bf Lemma}

\newtheorem{prop}[thm]{\bf Proposition}

\newtheorem{quest}[thm]{\bf Question}
\newtheorem{prob}[thm]{\bf Problem}

\theoremstyle{definition}
\newtheorem{defn}[thm]{\bf Definition}
\newtheorem{convent}[thm]{\bf Conventions}
\newtheorem{rem}[thm]{\bf Remark}
\newtheorem{ex}[thm]{\bf Example}

\title[Prime ideals and regular sequences]{Prime ideals and regular sequences of symmetric polynomials}
\author{Neeraj~Kumar} 
\address{Dipartimento di Matematica, Universit\'{a} di Genova\\
Via Dodecaneso 35, 16146 Genova, Italy}
\email{kumar@dima.unige.it}

\thanks{{\it Key words:} Regular sequence, Prime ideal, Symmetric polynomial.
\endgraf
{\it 2010 Mathematics Subject Classification:} Primary $13A15$, Secondary $14M10,\;12E05$.}

\begin{document}

\begin{abstract}
Let $S=\mathbb{C}[x_1,\dots,x_n]$ be a polynomial ring. Denote by $p_a$ the power sum 
symmetric polynomial $x_1^a+\cdots+x_n^a$. We consider the following two questions: 
Describe the subsets $A \subset \mathbb{N}$ such that the set of 
polynomials $p_{a}$ with $a \in A$ generate a prime ideal in $S$ or the set of 
polynomials $p_{a}$ with $a \in A$ is a regular sequence in $S$. We produce a large 
families of prime ideals by exploiting Serre's criterion for normality \cite[Theorem $18.15$]{DE} 
with the help of arithmetic considerations, vanishing sums of roots of 
unity \cite{Lam}. We also deduce several other results concerning 
regular sequences of symmetric polynomials. 
\end{abstract}

\maketitle

\section{Introduction}
Let $S=\mathbb{C}[x_1,\dots,x_n]$ be a polynomial ring. A sequence of elements 
$y_1,y_2,\dots,y_d$ in a ring $S$ is called \emph{regular sequence} on $S$ 
if the ideal $\langle y_1,y_2,\dots,y_d \rangle$ is proper and for each $i$, the image of $y_{i+1}$ is a 
nonzero divisor in $S/{ \langle y_1,\dots,y_i \rangle }$.

Following the notation of Macdonald \cite{I-G}, let $p_a,h_a$ and $e_a$ denote 
the power sum symmetric polynomial, complete symmetric polynomial, and the 
elementary symmetric polynomial of degree $a$ in $S$ respectively. Let $\mathbb{N}$ be 
the set of positive integers. For a given set $A \subset \mathbb{N}$, we denote by 
the set of power sum symmetric polynomials as $p_{A}=\{ p_a \; |\; a \in A \}$. 
In this paper, we discuss the following two questions:
\begin{quest}\label{question-2} Let $S=\mathbb{C}[x_1,\dots,x_n]$ be a polynomial ring. For 
which subsets $A \subset \mathbb{N}$, the ideal generated by the set of polynomials 
$p_{A}$ is a prime ideal in $S$.
\end{quest}
\begin{quest}\label{question-1} Let $S=\mathbb{C}[x_1,\dots,x_n]$ be a polynomial ring. For which 
subsets $A \subset \mathbb{N}$, the set of polynomials $p_{A}$ is a regular sequence in $S$. 
\end{quest}
Similarly, we ask these questions for the complete symmetric polynomials and the 
elementary symmetric polynomials. For obvious reasons, whenever $p_A$ is a regular 
sequence generating a prime ideal and $p_b \notin \langle p_{A} \rangle$, 
then $p_{A},p_b$ is a regular sequence in $S$ as well. More specifically, we will focus on 
Question \ref{question-2}.

These problems are fundamental in nature, and interesting, both in algebraic and 
geometric point of view. In a more general setting of Question \ref{question-2}, 
it is mentioned by Eisenbud 
that `` In general it is extremely difficult to prove that a given ideal of 
polynomial is prime'' \cite[Chapter $10$: pg.$241$]{DE}. 
The most powerful methods known for showing primeness of ideal are Hochster's 
method of \emph{``principal radical 
system''} \cite{Hos} and \emph{Serre's criterion for normality} \cite[Theorem $18.15$]{DE}. We will use the 
Serre's criterion for normality in this paper.  

The study of Question \ref{question-1} began in the paper \cite{CKW} in the dimention zero case by 
Conca, Krattenthaler, and Watanabe. The Question \ref{question-1} is highly non-trivial for $n\geq 3$. 
For $n=3$, a beautiful conjecture of Conca, Krattenthaler, and Watanabe states that given a positive 
integers $a<b<c$ with $\gcd(a,b,c)=1$, $p_a,p_b,p_c$ form a regular sequence in 
$\mathbb{C}[x_1,x_2,x_3]$ if and only if $abc \equiv 0 ( \mod 6)$, see  \cite[Conjecture $2.15$]{CKW}. 
The necessary condition follows from \cite[Lemma $2.8$]{CKW}. For partial evidence in support of sufficient 
condition, see \cite[Theorem $2.11$]{CKW}. Similarly the authors, also formulated a conjecture, 
when three complete symmetric polynomials form a regular sequence in $\mathbb{C}[x_1,x_2,x_3]$, 
see  \cite[Conjecture $2.17$]{CKW}. In a joint paper with Martino \cite{KM}, we could provide 
evidence for these conjectures 
by proving it in special instances. Then we employed the technique of Serre's 
criterion to show the primeness of an ideal. For instance, we have shown that the 
ideal $I=\langle p_{a}, p_{a+1},\dots, p_{a+ m-1} \rangle $ is prime in $S$
if $m < {n-1}$,  see \cite[Theorem $3.3$]{KM}. 
We have also shown that the ideal $ I=\langle p_1,p_{2m} \rangle $, where $m \in \mathbb{N}$, is prime 
in $\mathbb{C}[x_1,\dots,x_4]$, see \cite[Proposition $4.1$]{KM}. In this way, we succeeded to 
give more families of regular sequences. With the help of 
Computer calculations, we proposed, Conjecture $4.5$ and Conjecture $4.6$ in \cite{KM}. 
One of the main result of this paper, Theorem \ref{main-thm-prime-ideal-any-n}, partially 
answers the Conjecture $4.6$ in \cite{KM}.

In this paper, we have managed to produce families of prime ideals by exploiting Serre's criterion 
with the help of arithmetic considerations, vanishing sums of roots of unity \cite{Lam}. 
The main results of the paper are the following: 
\begin{itemize}
\item[(i)] Let $S=\mathbb{C}[x_1,x_2,\dots,x_n]$ 
be a polynomial ring with $n \geq 4$. Let $I=\langle p_{a},p_{b} \rangle$, where $a,b \in \mathbb{N}$. 
Let $b-a=n_0$. Suppose $q_1$ is the smallest prime factor 
in the factorization of $n_0$. If $q_1 > \max \{n, a\}$, then $I$ is a prime 
ideal in $S$. 
\item[(ii)] Let $S=\mathbb{C}[x_1,x_2,\dots,x_n]$ be a polynomial ring with $n \geq 3$. Let 
$a \in \mathbb{N}$ and $m < n-1$. Let $I=\langle p_{a},p_{2a},\dots,p_{ma} \rangle $. 
Then $I$ is a prime ideal in $S$. 
\end{itemize}
Section $2$ contains preliminary results. In Section $3$, we discuss the problem 
of whether two power sum symmetric polynomials generate a prime ideal in $S$ for $n \geq 4$. We answer this to  
certain extent purely in terms of arithmetic conditions of the degree of polynomials and the number of 
indeterminates, see Theorem \ref{main-thm-prime-ideal-any-n}. 
Let $I= \; \langle p_{a},p_{2a},\dots,p_{ma} \rangle$, where $a \in \mathbb{N}$, and $m < n-1$. 
We show that $I$ is a prime ideal in $S$ for all $n \geq 3$, see Theorem \ref{p-a-2a-upto ma-prime ideal}. 
For $n \geq 3$, we show that $ \frac{\partial h_a}{\partial x_1},\cdots, \frac{\partial h_a}{\partial x_n}$ form a 
regular sequence in $S$ for all $a \geq 2$. We also show that any two complete symmetric 
polynomial form a regular sequence in $S$ for all $n \geq 3$. Similar results also hold for the power sum 
and elementary symmetric polynomials, see Lemma \ref{partial-derivates-reg-seq} and 
Proposition \ref{ha-hb}. Computer calculations in CoCoA 
\cite{CoCoA} suggest that $I=\langle h_{1},h_{2m}\rangle$, where $m \in \mathbb{N}$, 
should be a prime ideal in $\mathbb{C}[x_1,\dots,x_4]$. It is obvious for $m=1$. We provide 
evidence for $m=2$ in the Example \ref{h-1-4}.
\section{Generalities and preliminary results}

\medskip

Let $S=\mathbb{C}[x_1,\dots,x_n]$ be a polynomial ring. Let $p_a,\;h_a$ and $e_a$ be the 
power sum symmetric polynomial, complete symmetric polynomial, and the 
elementary symmetric polynomial of degree $a$ in $S$ respectively, that is,
\[
 \begin{split}
 p_a(x_1,x_2,\dots,x_n):=& \sum_{i=1}^{n}x_i^a, \\
 h_a(x_1,x_2,\dots,x_n):=& \sum_{1 \leq i_{1} \leq i_{2} \leq \cdots \leq i_a \leq n} x_{i_1}x_{i_2}\cdots x_{i_a},\\
 e_a(x_1,x_2,\dots,x_n):=& \sum_{1 \leq i_{1} < i_{2} < \cdots < i_a \leq n} x_{i_1}x_{i_2}\cdots x_{i_a}.
 \end{split}
\]
For instance, for $n=3$ and $a=2$, one has 
\[
\begin{split}
 p_2(x_1,x_2,x_3)=& x_1^2+x_2^2+x_3^2, \\
 h_2(x_1,x_2,x_3)=& x_1^2+x_2^2+x_3^2+x_1x_2+x_1x_3+x_2x_3,\\
 e_2(x_1,x_2,x_3)=& x_1x_2+x_1x_3+x_2x_3.
 \end{split}
\]
For these symmetric polynomials, we have the following \emph{Newton's formula}, 
see \cite[Equation's $2.6^{\prime},2.11, 2.11^{\prime}$ respectively]{I-G}: 
\begin{eqnarray}\label{Newton's formula-2}
  \sum_{i=0}^{n} (-1)^{i}e_i h_{n-i} =0 \text{ for all $n \geq 1,$ }
\end{eqnarray}
\begin{eqnarray}\label{h-p relation}
 ah_a= \sum_{i=1}^{a} p_i h_{a-i}  \text{ for all $a \geq 1,$ }
\end{eqnarray}
\begin{eqnarray}\label{Newton's formula-1}
 ne_n=\sum_{i=1}^{n} (-1)^{i-1}e_{n-i} p_{i} \text{ for all $n \geq 1$. } 
\end{eqnarray}
We will use the following lemma to prove the smoothness of symmetric polynomials 
$h_a,\;e_a$, and $p_a$ in the Lemma \ref{p-h-a-n variable}. We will also use 
Lemma \ref{h-poly} in the Example \ref{h-1-4}. 

\begin{lem}\label{h-poly} {\rm (Technical Lemma)} \\[1mm]
Let $S=\mathbb{C}[x_1,\dots,x_n]$ be a polynomial ring. Then for the symmetric polynomials 
$h_a,\;e_a$, and $p_a$, one has the following:
 \begin{itemize}
\item[(i)]  $\frac{\partial h_a}{\partial x_i} = h_{a-1} + x_i \frac{\partial h_{a-1}}{\partial x_i}$ and  
$\sum_{i=1}^{n}\frac{\partial h_a}{\partial x_i}=(n+a-1)h_{a-1}$.
\item[(ii)] $\frac{\partial e_a}{\partial x_i} = e_{a-1} - x_i \frac{\partial e_{a-1}}{\partial x_i}$ and 
 $\sum_{i=1}^{n}\frac{\partial e_a}{\partial x_i}=(n-a+1)e_{a-1}$.
\item [(iii)] $\frac{\partial p_a}{\partial x_i} = {a}x_i^{a-1}$ and 
$\sum_{i=1}^{n}\frac{\partial p_a}{\partial x_i}= ap_{a-1}$.
\end{itemize}
\end{lem}
\begin{proof} We can write $h_a=  x_i h_{a-1} + g$ for some polynomial $g$ not involving $x_i$. 
Taking partial derivative of $h_a$ w.r.t. $x_i$, one obtains 
 \begin{eqnarray}\label{h-derivative}
 \frac{\partial h_a}{\partial x_i} = h_{a-1} + x_i \frac{\partial h_{a-1}}{\partial x_i}.
\end{eqnarray}
By Euler's formula, we have $\sum_{i=1}^{n} x_i \frac{\partial h_{a-1}}{\partial x_i}=(a-1)h_{a-1}.$
Thus, we conclude that
\begin{eqnarray*}\label{sum-h-derivative}
 \sum_{i=1}^{n}\frac{\partial h_a}{\partial x_i}=(n+a-1)h_{a-1}.
\end{eqnarray*}
Taking partial derivative of $e_a$ w.r.t. $x_i$, one obtains 
\begin{eqnarray}\label{e-derivative}
 \frac{\partial e_a}{\partial x_i} = e_{a-1} - x_i \frac{\partial e_{a-1}}{\partial x_i}.
\end{eqnarray}
Proceeding as before, we conclude that
\begin{eqnarray*}\label{sum-e-derivative}
 \sum_{i=1}^{n}\frac{\partial e_a}{\partial x_i}=(n-a+1)e_{a-1}.
\end{eqnarray*}
The claim of (iii) is obvious.
\end{proof}

For a given natural number $m$, consider the $m$-th roots of unity in the field of 
complex number $\mathbb{C}$. We may ask ourself for which natural numbers 
$n$ and $k$, do there exist $m$-th roots of unity $\alpha_1,\dots,\alpha_n \in \mathbb{C}$ such that 
$\alpha_1^k+\alpha_2^k+\dots+\alpha_n^k=0$? We define such an equation to be 
\emph{Vanishing sum of $k$-th power of $m$-th roots of unity of weight $n$}. 
For $k=1$, such an equation is said to be a \emph{vanishing sum of $m$-th 
roots of unity of weight $n$}. For instance, for $m=10$ and $k=1$, the set of $n$'s 
is $\{0,2,4,5,6,7,8,9,10,\dots \}$. For a given $m$ and $k$, let $W(m,k)$ be the set of weights $n$ for which 
there exists a vanishing sum $\alpha_1^k+\alpha_2^k+\dots+\alpha_n^k=0$, where 
$\alpha_i$ is a $m$-th roots of unity. If $m$ has a prime 
factorization of the form $q_1^{a_1}q_2^{a_2}\dots q_r^{a_r}$, then by the theorem of 
Lam and Leung \cite[Theorem $5.2$]{Lam}, the weight set $W(m,1)$ is exactly given 
by $\mathbb{N}q_1+\dots+\mathbb{N}q_r$. Poonen and Rubinstein \cite{Poonen-Rub} have 
classified all minimal vanishing sums $\alpha_1+ \cdots+\alpha_n=0$ of weight $n \leq 12$. 
For similar treatment by Mann, one may also see \cite{Mann}.

\begin{rem}\label{lam-remark}
 Note that by \cite[Theorem $5.2$]{Lam}, for a given $m \in \mathbb{N}$, $W(m,1)$ depends only on the prime 
 divisors of $m$, and not on the multiplicity to which they occur in the 
 factorization of $m$. The theorem also shows that any (non-empty) vanishing 
 sum of $m$-th roots of unity must have weight $\geq q_1$, where $q_1$ is 
 the smallest prime divisor of $m$.
\end{rem}

\begin{lem}\label{imp-lemma}
Let $m \in \mathbb{N}$ and $\alpha_1,\dots,\alpha_n$ be a subset of $m$-th 
roots of unity in $\mathbb{C}$. Suppose that $q_1$ is the smallest prime factor 
in the factorization of $m$. Let $k \in \mathbb{N}$ be any natural number such 
that $q_1 > \max \{n, k\}$. Then one has $\alpha_1^k+\alpha_2^k+\cdots+\alpha_n^k \neq 0$.
\end{lem}
\begin{proof}
 The claim follows from Remark \ref{lam-remark}.
\end{proof}
\medskip

\section{Prime ideals and regular sequences}
\medskip

We begin this section with a useful lemma, which states that 
all the partial derivatives of a complete symmetric polynomial form a regular sequence in the polynomial 
ring, see Lemma \ref{partial-derivates-reg-seq}. Then we use Lemma \ref{partial-derivates-reg-seq} to show 
that the complete symmetric polynomials and their partial derivatives are smooth polynomials, 
see Lemma \ref{p-h-a-n variable} and Lemma \ref{partial-derivates-h-smooth}. Then we recall the 
irreducibility of Schur polynomial in the polynomial ring, see \cite[Theorem $3.1$]{DZ-2}. In a special 
case, we discuss the smoothness of Schur polynomial, see Example \ref{Schur-poly-irre}. 
Then we discuss the main results of this paper, Theorem \ref{main-thm-prime-ideal-any-n} and 
Theorem \ref{p-a-2a-upto ma-prime ideal} respectively. We also discuss the case of two complete symmetric polynomials 
generating a prime ideal in the polynomial ring. 

\begin{lem}\label{partial-derivates-reg-seq}
 Let $n \in \mathbb{N}$ with $n \geq 3$. Let $S=\mathbb{C}[x_1,x_2,\dots,x_n]$ be a polynomial ring. 
 Let $a \geq 2$. Then the following holds:
 \begin{itemize}
\item[(i)] $ \frac{\partial h_a}{\partial x_1},\cdots, \frac{\partial h_a}{\partial x_n}$ form 
a regular sequence in $S$.
\item[(ii)] $ \frac{\partial p_a}{\partial x_1},\cdots, \frac{\partial p_a}{\partial x_n}$ form 
a regular sequence in $S$.
\item[(iii)] $ \frac{\partial e_a}{\partial x_1},\cdots, \frac{\partial e_a}{\partial x_n}$ form 
a regular sequence in $S$ for all $a < n$.
\end{itemize}
 \end{lem}
\begin{proof}
Let the ideal generated by all the partial derivatives of $h_a$ be $J_a$, that is,
 \begin{eqnarray}
  J_a= \langle \frac{\partial h_a}{\partial x_1},\cdots, \frac{\partial h_a}{\partial x_n}   \rangle.
 \end{eqnarray}
 We have $ah_a= \sum x_i\frac{\partial h_a}{\partial x_i}$. Thus, clearly $h_a \in J_a$. 
 By Lemma \ref{h-poly} (i), we have 
 \[
 (n+a-1)h_{a-1}=  \sum_{i=1}^{n}\frac{\partial h_a}{\partial x_i} \; \in J_a.
  \]
 That is, $h_{a-1} \in J_a$. By Lemma \ref{h-poly} (i), we have 
 \[
   \frac{\partial h_{a+1}}{\partial x_i}=h_{a}+x_i\frac{\partial h_a}{\partial x_i} \; \in J_a.
 \]
Thus, we observe that $J_{a+1} \subset J_a$. Moreover, $h_{a+1} \in J_{a+1}$. Proceeding similarly, we have a 
chain of containment of ideals
\[
 J_{a+n-2} \subset \cdots \subset J_{a+1} \subset J_{a}.
\]
Therefore, we have $h_{a-1}, h_a,h_{a+1},\cdots \subset h_{a+n-2} \in J_a$. 
Recall that by \cite[Proposition $2.9$]{CKW}, any $n$ consecutive complete symmetric polynomials 
$h_{a-1}, h_{a},\dots,h_{a+n-2}$ form a regular sequence in $S$ for all $a \geq 2$. 
Thus, we conclude that $\text{ht}(J_a)=n$, moreover $J_a$ is a complete intersection ideal in $S$. 
The claim (ii) is obvious.  Let the ideal generated by all the partial derivatives of $e_a$ be $E_a$. 
The proof of (iii) is similar to above, except the fact that this time, we choose carefully $a$'s 
such that $e_a \in E_a$. The reason for this is that $e_a=0$ for all $a >n$. We want to use the 
fact that $e_1,e_2,\dots,e_n \in E_a$. As we know that $e_1,\cdots,e_n$ form a regular sequence in $S$. 
\end{proof}

In the following lemma, we discuss the smoothness of the symmetric polynomials 
$h_a,\;e_a$ and $p_a$. 

\begin{lem}\label{p-h-a-n variable}
 Let $n \in \mathbb{N}$ with $n \geq 3$. Let $S=\mathbb{C}[x_1,x_2,\dots,x_n]$ be a polynomial ring.
 Then the following holds:
 \begin{itemize}
\item[(i)] $h_a$ is smooth, hence an irreducible element in $S$ for all $a \geq 1$.
\item[(ii)] $e_a$ is smooth, hence an irreducible element in $S$ for all $1 \leq a \leq n-1$.
\item[(iii)] $p_a$ is smooth, hence an irreducible element in $S$ for all $a \geq 1$.
\end{itemize}
 \end{lem}
\begin{proof} If $a=1$, the claims are obvious. We will give an elementary proof of $(i)$. 
The proof of (ii) and (iii) are similar.\\
{\bf{Proof of $(i)$:}} If $h_a= f \cdot g$ with $f$ and $g$ non constant polynomial, 
then $f$ and $g$ have to be homogeneous. 
By Bezout theorem, the hypersurfaces $f=0$ and $g=0$ intersects in the projective space 
$\mathbb{P}^{n-1}$, since $n \geq 3$. This gives a singular point on the hypersurface $h_a=0$. 
So, it suffices to prove that $h_a,\frac{\partial h_{a}}{\partial x_1},\dots,\frac{\partial h_{a}}{\partial x_n}$ 
have no common zero in $\mathbb{C}^n-\{0\}$. This claim follows from Lemma \ref{partial-derivates-reg-seq} (i).
\end{proof}

In the following lemma, we discuss the smoothness of the partial derivatives of the complete symmetric polynomials. 

\begin{lem}\label{partial-derivates-h-smooth}
Let $n \in \mathbb{N}$ with $n \geq 3$. Let $S=\mathbb{C}[x_1,x_2,\dots,x_n]$ be a polynomial ring. 
 Let $a \geq 3$. Then $ \frac{\partial h_a}{\partial x_i}$ is smooth, and hence irreducible in $S$.
\end{lem}
\begin{proof}
It is enough to show that $ \frac{\partial h_a}{\partial x_1}$ is smooth. Similar to the 
proof given in Lemma \ref{p-h-a-n variable}, it suffices to show that all its partial derivatives 
$ \frac{\partial^2 h_a}{\partial x_1 \partial x_1}, \frac{\partial^2 h_a}{\partial x_1 \partial x_2},\dots, 
\frac{\partial^2 h_a}{\partial x_1 \partial x_n}$ have no common zero in $\mathbb{C}^n-\{0\}$. We set few notations 
for convenience, for instance, let $g_{a,i}=\frac{\partial h_a}{\partial x_i}$. Let the ideal generated by 
all the partial derivatives of $g_{a,1}$ be $I_a$:
\[
 I_a=\langle   \frac{\partial g_{a,1}}{\partial x_1}, \frac{\partial g_{a,1}}{\partial x_2}\cdots, \frac{\partial g_{a,1}}{\partial x_n}  \rangle. 
\]
Also note that $ \frac{\partial g_{a,j}}{\partial x_i}=\frac{\partial g_{a,i}}{\partial x_j}$. With this observation, 
one obtains 
\[
 I_a=\langle   \frac{\partial g_{a,1}}{\partial x_1}, \frac{\partial g_{a,2}}{\partial x_1}\cdots, 
 \frac{\partial g_{a,n}}{\partial x_1}  \rangle.
\]
Taking partial derivatives of $g_{a,i}$ w.r.t. $x_1$, for $i=1,\dots,n$,
\[
 \begin{split}
 \frac{\partial g_{a,1}}{\partial x_1}=& \frac{\partial h_{a-1}}{\partial x_1}+ 
 x_1 \frac{\partial^2 h_a}{\partial x_1 \partial x_1} + \frac{\partial h_{a-1}}{\partial x_i},\\
 \frac{\partial g_{a,2}}{\partial x_1}=& \frac{\partial h_{a-1}}{\partial x_1}+ 
 x_2 \frac{\partial^2 h_a}{\partial x_2 \partial x_1}\\
 \vdots& \\
  \frac{\partial g_{a,n}}{\partial x_1}=& \frac{\partial h_{a-1}}{\partial x_1}+ 
 x_n \frac{\partial^2 h_a}{\partial x_n \partial x_1}
 \end{split}
\]
Then summing up, we get
\[
 \begin{split}
 \sum_{i=1}^{n} \frac{\partial g_{a,i}}{\partial x_1}=& (n+1)\frac{\partial h_{a-1}}{\partial x_1} +  \sum_{i=1}^{n} 
 x_i \frac{\partial }{\partial x_i}  ( \frac{\partial h_{a-1}}{\partial x_1}  ) \\
 =& (\sharp)\frac{\partial h_{a-1}}{\partial x_1},
  \end{split}
 \]
for some integer number $\sharp$, which is irrelevant. Thus, we conclude that 
$\frac{\partial h_{a-1}}{\partial x_1} \in I_a$. Also $g_{a,1} \in I_a$. 
We also know that $g_{a,1}=\frac{\partial h_a}{\partial x_1}=h_{a-1}+x_1 \frac{\partial h_{a-1}}{\partial x_1}$. 
Thus, we get $h_{a-1} \in I_a$. Now consider the ideal 
\[
 I_{a+1}=\langle   \frac{\partial g_{a+1,1}}{\partial x_1}, \frac{\partial g_{a+1,1}}{\partial x_2}\cdots, 
 \frac{\partial g_{a+1,1}}{\partial x_n}  \rangle. 
\]
Proceeding similarly, we obtain
\[
 \begin{split}
 \sum_{i=1}^{n} \frac{\partial g_{a+1,i}}{\partial x_1}=& (n+1)\frac{\partial h_{a}}{\partial x_1} +  \sum_{i=1}^{n} 
 x_i \frac{\partial }{\partial x_i}  ( \frac{\partial h_{a}}{\partial x_1}  ) \\
 =& (\sharp_1)\frac{\partial h_{a}}{\partial x_1},
  \end{split}
\]
for some integer number $\sharp_1$, which is irrelevant. Thus, we conclude that 
$\frac{\partial h_{a}}{\partial x_1} \in I_{a+1}$. Also $g_{a+1,1} \in I_{a+1}$. 
We also know that $g_{a+1,1}=\frac{\partial h_{a+1}}{\partial x_1}=h_{a}+x_1 \frac{\partial h_{a}}{\partial x_1}$. 
Thus, we get $h_{a} \in I_{a+1}$. Also note that $\frac{\partial g_{a+1,1}}{\partial x_i} \in I_a $ for all $i$. 
That is, $I_{a+1} \subset I_a$. Proceeding in a similar way, we obtain the following relations
\[
 I_{a+j+1} \subset I_{a+j} \text{, and } h_{a+j-1} \in I_{a+j}  
\]
for all $j \geq 0$. Thus, using similar argument as in the Lemma \ref{partial-derivates-reg-seq}, we 
 conclude that $\text{ht}(I_a)=n$, moreover $I_a$ is a complete intersection ideal in $S$. 
Thus the claim follows.
\end{proof}

Following Macdonald \cite{I-G}, the \emph{Schur polynomial} is defined as
\[
 s_{\lambda}=s_{\lambda}(x_1,\dots,x_n) =
 \frac{\dett (x_i^{\lambda_j+n-j})_{1 \leq i,j \leq n} }{ \dett (x_i^{n-j})_{1 \leq i < j \leq n} },
\]
where $\lambda= ( \lambda_1,\lambda_2 , \dots, \lambda_n)$ is the partition of non-negative 
integers with $\lambda_i \geq \lambda_{i+1}$ for $i=1,\dots,{n-1}$.
For Schur polynomial, we have the following relation, see \cite[Equation $3.4$]{I-G}: 
\begin{eqnarray}\label{s and h relation}
s_{\lambda}=\dett ( h_{{\lambda_i}-i+j})_{1 \leq i,j \leq n} 
\end{eqnarray}
where $n \geq l(\lambda)$. 

\begin{rem}
Irreducibility of Schur polynomial is discussed by Dvornicich and Zannier in \cite{DZ-2}. Let $n \geq 3$. 
For a given partition $\lambda= ( \lambda_1,\lambda_2 , \dots, \lambda_n)$, where $\lambda_1
> \lambda_2 > \cdots > \lambda_n $ with $\lambda_n=0$ and 
$\gcd(\lambda_1,\lambda_2 , \dots, \lambda_{n-1})=1$, then the Schur polynomial $s_{\lambda}(x_1,\dots,x_n)$ is 
irreducible in $S=\mathbb{C}[x_1,x_2,\dots,x_n]$, see \cite[Theorem $3.1$]{DZ-2}. 
Recall that $h_a$ and $e_a$ are special forms of a Schur polynomial. 
Simply note that the irreducibility of $h_a$ and $e_a$ does 
not follow from \cite[Theorem $3.1$]{DZ-2}. In the Lemma \ref{p-h-a-n variable}, we not only show the 
irreducibility of these polynomials, but also the smoothness. The obtained results partially extend the domain 
of partition for the irreducibility of Schur polynomial.
\end{rem}

Assuming that the Schur polynomial $s_{\lambda}(x_1,\dots,x_n)$ is irreducible in $S$. 
One may ask, whether is it true that the Schur polynomial is also smooth? The answer is positive in the 
case of $h_a$ and $e_a$. Computational evidence shows that the answer is negative in general. However, in a 
special case, when the partition $\lambda$ is of the form $(\lambda_1,1,0)$ for $\lambda_1 \geq 2$, 
then the Schur polynomial $s_{\lambda}(x_1,x_2,x_3)$ turns out to be smooth in $\mathbb{C}[x_1,x_2,x_3]$. 
We discuss the proof of this in the following example:
\begin{ex}\label{Schur-poly-irre}
Let $S=\mathbb{C}[x_1,x_2,x_3]$ be a polynomial ring. Let $\lambda= (\lambda_1,1,0)$ 
be the partition. Then the Schur polynomial $s_{\lambda}(x_1,x_2,x_3)$ is smooth
in $S$ for all $\lambda_1 \geq 2$.
\end{ex}
\begin{proof}
 By $(\ref{s and h relation})$, we have 
\[
 s_{\lambda}=s_{(\lambda_1,1,0)}=h_1h_{\lambda_1}-h_{\lambda_1+1}.
\]
Similar to the proof given in Lemma \ref{p-h-a-n variable}, it suffices to show that all its partial derivatives 
$\frac{\partial s_{\lambda}}{\partial x_1},\frac{\partial s_{\lambda}}{\partial x_2},
\frac{\partial s_{\lambda}}{\partial x_3} $ have no common zero in $\mathbb{C}^3-\{0\}$. 
Taking partial derivatives of $s_{\lambda}$ w.r.t. $x_i$, we get
\[
\begin{split}
 \frac{\partial s_{\lambda}}{\partial x_i}= & h_{\lambda_1} + h_1 \frac{\partial h_{\lambda_1}}{\partial x_i}-
 \frac{\partial h_{\lambda_1+1}}{\partial x_i}\\
 =& (h_1-x_i) \frac{\partial h_{\lambda_1}}{\partial x_i} \text{  for all $i=1,2,3$.}
\end{split}
 \]
We see that $h_1\neq x_i$, unless $x_i=0$ for all $i$. Thus the common zero of $ 
\frac{\partial h_{\lambda_1}}{\partial x_1}, 
 \frac{\partial h_{\lambda_1}}{\partial x_2},  \frac{\partial h_{\lambda_1}}{\partial x_3} $
is also a common zero of $\frac{\partial s_{\lambda}}{\partial x_1},\frac{\partial s_{\lambda}}{\partial x_2},
\frac{\partial s_{\lambda}}{\partial x_3} $.  
By Lemma \ref{partial-derivates-reg-seq}, we conclude that
$ \langle \frac{\partial s_{\lambda}}{\partial x_1}, \frac{\partial s_{\lambda}}{\partial x_2}, 
 \frac{\partial s_{\lambda}}{\partial x_3} \rangle $ is a complete intersection ideal in $S$. 
 Thus the claim follows.
\end{proof}

We record the following convention, which we will follow from here onwards throughout this paper:
\begin{convent} When we list the symmetric polynomials $f_{i_1},f_{i_2},\dots, f_{i_k}$ with respective 
degrees $\deg (f_{i_j})=i_j$, we always assume that $i_1<i_2<\cdots <i_k$, unless otherwise specified. 
\end{convent}

In the following proposition, we will see that any two complete symmetric polynomials 
always form a regular sequence in $S=\mathbb{C}[x_1,\dots,x_n]$ for $n \geq 3$. 
Similar results also hold for the power sum and elementary symmetric polynomials.

\begin{prop}\label{ha-hb}
Let $n \in \mathbb{N}$ with $n \geq 3$. Let $S=\mathbb{C}[x_1,x_2,\dots,x_n]$ be a polynomial ring. 
Then the following holds:
\begin{itemize}
 \item [(i)] $h_a,h_b$ form a regular sequence.
 \item [(ii)] $e_a,e_b$ form a regular sequence for all $1 \leq a <b \leq n-1$.
 \item [(iii)] $p_a,p_b$ form a regular sequence.
\end{itemize}
\end{prop}
\begin{proof} We will prove (i). By Lemma \ref{p-h-a-n variable}, $h_a$ is an irreducible polynomial in $S$. 
Hence $S/{ \langle h_a \rangle }$ is a domain. Now $h_b$ being an irreducible polynomial in $S$, 
can not be factored into lower degree complete symmetric polynomials $h_a$. 
So, $h_b$ is a nonzero divisor in $S/{\langle h_a \rangle }$ for $b > a$. Hence $h_a,h_b$ form a regular sequence in $S$. 
Proof of (ii) and (iii) are similar.
\end{proof}

\medskip

We have seen that any two power sum polynomials $p_a,p_b$ form a regular sequence
in $S=\mathbb{C}[x_1,x_2,\dots,x_n]$ for $n \geq 3$. We would like to know when two 
power sum polynomials generate a prime ideal in $S$ for $n \geq 4$. In a special case, some 
answers are known due to \cite[Theorem $4.3$ and Proposition $4.3$]{KM}. 
In the following theorem, we will answer this to certain extent, when two power sum 
polynomials generate a prime ideal in $S$ for $n \geq 4$, purely in terms of arithmetic 
conditions of the degree of polynomials and the number of indeterminates.

\begin{thm}\label{main-thm-prime-ideal-any-n} Let $S=\mathbb{C}[x_1,x_2,\dots,x_n]$ 
be a polynomial ring with $n \geq 4$. Let $I=\langle p_{a},p_{b} \rangle$, 
where $a,b \in \mathbb{N}$. Let $b-a=n_0$. Suppose $q_1$ is the smallest prime factor 
in the factorization of $n_0$. If $q_1 > \max \{n, a\}$, then $I$ is a prime 
ideal in $S$. Moreover $p_{a},p_{b},p_c$ form a regular sequence in $S$ for 
all $p_c \notin \langle p_{a},p_{b} \rangle$. 
\end{thm}
\begin{proof}
If $n_0=1$, then $I$ is a prime ideal in $S$ follows from \cite[Theorem $4.3$]{KM}. 
Thus, we assume that $n_0 \geq 2$. Let $R={S}/{I}$. We compute the Jacobian of 
$I$ up to scalar (We can ignore the coefficients, since we are in the field of 
characteristic zero.), say Jacobian is $J$:
\[ J=
\begin{pmatrix}  x_1^{a-1} &  x_2^{a-1} & \dots & x_n^{a-1}\\  
                 x_1^{b-1} &  x_2^{b-1} & \dots &  x_n^{b-1}\\
 \end{pmatrix}.
\]
Let $J^{\prime}= I_2(J)$, denotes the ideal generated by $2 \times 2$ minors 
of Jacobian. Also $\text{ht}(I)=2$, since $I$ is generated by a regular 
sequence of length $2$. The determinants of $2 \times 2$ minors of the 
Jacobian can be written as 
\[
 J^{\prime}=\langle \; x_j^{a-1}x_i^{a-1}(x_j^{b-a}-x_i^{b-a})   \;\rangle \text{ for $1 \leq i<j \leq n$. }
\]
Thus, we have
\[ 
I+J^{\prime}=\langle \; p_{a}, p_{b},\;x_j^{a-1}x_i^{a-1}(x_j^{b-a}-x_i^{b-a}) \; \rangle 
\text{ for $1 \leq i<j \leq n$. }
\]
Claim: $\sqrt{I+J^{\prime}}=( x_1,x_2,\dots,x_n).$ \\
Suppose not, that is, there exists $w =(w_1,w_2,\dots,w_n) \in \mathbb{P}^{n-1}$ 
with $w \in Z(I+J^{\prime})$.
 Since $w$ is in $\mathbb{P}^{n-1}$, we may assume $w=(1,y_1,y_2,\dots,y_{n-1})$. 
 As $w \in Z(I+J^{\prime})$, we have that $y_i^{b-a}=y_i^{n_0}=1$ for all $i=1,\dots {n-1}$, 
 moreover $w$ also satisfies $p_{a}, p_{b}$. Therefore we have
\[
 1+y_1^a+y_2^a+\cdots+y_{n-1}^a=0 \text{ and } 1+y_1^b+y_2^b+\cdots+y_{n-1}^b=0.
\]
Both the equations reduce to the existence of solution of $1+y_1^a+y_2^a+\cdots+y_{n-1}^a=0$. 
We use the fact that all the $y_i$'s are $n_0$-th roots of unity, say $1,\zeta_1,\dots,\zeta_{n_0 -1}$. 
Suppose $q_1$ is the smallest prime factor in the factorization of $n_0$. 
If $q_1 > \max \{n, a\}$, then it follows from Lemma \ref{imp-lemma} that 
$1+y_1^a+y_2^a+\cdots+y_{n-1}^a \neq 0$. So, the only possible solution has 
to be the trivial solution. 
Hence the claim is proved. 

Thus $\text{ht}(I+J^{\prime})=n$ and $\dim {S}/{(I+J^{\prime})}=0$. The co-dimension of 
$J^{\prime}$ in $S$ is $n-2$. By  \cite[Theorem $18.15$]{DE}, $R$ is a product of normal 
domain, since $n \geq 4$. Thus, we can write $R=R_1\times \cdots \times R_k$. 
Since $R$ is a standard graded $\mathbb{C}$-algebra with $R_0=\mathbb{C}$, 
also $R_0=(R_1)_0\times \cdots \times (R_k)_0=\mathbb{C}^k$. Hence $k=1$. 
Thus $R$ is a normal domain and $I$ is a prime ideal in $S$.
\end{proof}

\begin{rem}\label{exception-pa-p2a}
Let $p_1,p_2,p_5 \in \mathbb{C}[x_1,\dots,x_4]$. By Newton's formula $(\ref{Newton's formula-1})$, 
we observe that $p_5\equiv 0 \mod(p_1,p_2)$. Hence $p_5 \in \langle p_1,p_2 \rangle $. Replacing $x_i$ by $x_i^d$, we 
may also conclude that $p_{5d} \in \langle p_{d},p_{2d} \rangle $. Thus in the hypothesis of 
Theorem \ref{main-thm-prime-ideal-any-n}, we need the 
condition $p_c \notin \langle p_a,p_b \rangle $. 
\end{rem}

\begin{rem}
Computer calculations in CoCoA \cite{CoCoA} suggest that whenever 
$I=\langle p_{a},p_{b}\rangle$ is 
a prime ideal in $\mathbb{C}[x_1,x_2,x_3,x_4]$, then $p_{a},p_{b},p_c$ 
form a regular sequence, except $p_{a},p_{2a},p_{5a}$. It is clear from 
Remark \ref{exception-pa-p2a} that $p_{a},p_{2a},p_{5a}$ do not form a 
regular sequence. If this computational claim can be answered, then it will 
prove \cite[Conjecture $4.5$]{KM} to certain extent.
\end{rem}

\begin{prop}\label{generate more prime ideal}
Let $I$ and $J$ be the prime ideal in $K[x_1,\dots,x_n]$ and $K[y_1,\dots,y_m]$ 
respectively, where $K$ is an algebraically closed field. Let $(I,J)$ be the ideal 
generated by elements of $I$ and $J$ in $K[x_1,\dots,x_n,y_1,\dots,y_m]$. Then 
$(I,J)$ is a prime ideal in $K[x_1,\dots,x_n,y_1,\dots,y_m]$.
\end{prop}
\begin{proof}
 There is a standard isomorphism
 \[
  K[x_1,\dots,x_n]/I \tensor K[y_1,\dots,y_m]/J \cong 
  K[x_1,\dots,x_n,y_1,\dots,y_m]/(I,J)
 \]
by sending $f \tensor g \mapsto fg$. We see that both $K[x_1,\dots,x_n]/I$ and 
$K[y_1,\dots,y_m]/J$ are integral domains as well as $K$-algebras. By \cite[Proposition $4.15(b)$]{Milne}, 
the tensor product of $K[x_1,\dots,x_n]/I$ and $K[y_1,\dots,y_m]/J$ is also an integral domain, since 
$K$ is algebraically closed field. Hence the claim follows. 
\end{proof}

\begin{rem}
 Note that the goal of Proposition \ref{generate more prime ideal} is to generate more families of prime 
 ideals from given prime ideals.
\end{rem}
 
By \cite[Proposition $2.9$]{CKW}, we know that $p_1,p_2,\dots,p_n$ form a regular sequence 
in $S=\mathbb{C}[x_1,x_2,\dots,x_n]$. We also know that a subset of a 
regular sequence is again a regular sequence. Thus 
$p_1,p_2,\dots,p_m$ also form a regular sequence for all $m <n$. Then by \cite[Lemma $2.2$]{CKW}, 
we conclude that $p_{a},p_{2a},\dots,p_{ma}$ also form a regular sequence. Let 
$I= \langle p_{a},p_{2a},\dots,p_{ma} \rangle$, where $a \in \mathbb{N}$, and $m < n-1$. 
Let $R={S}/{I}$. Then $R$ is a Cohen-Macaulay ring. In the following theorem, we will show 
that $I$ is a prime ideal in $S$. We show this by proving that $R$ is a normal domain using Serre's 
criterion for normality.  
 
\begin{thm}\label{p-a-2a-upto ma-prime ideal} Let $S=\mathbb{C}[x_1,x_2,\dots,x_n]$ be a 
polynomial ring with $n \geq 3$. Let $a \in \mathbb{N}$. Let $I=\langle p_{a},p_{2a},\dots,p_{ma} \rangle $, 
where $m < n-1$. Then $I$ is a prime ideal in $S$. 
\end{thm}
\begin{proof}
For $a=1$, it follows from \cite[Proposition $4.3$]{KM}. Assume $a>1$. Let $R={S}/{I}$. 
We compute the Jacobian of $I$ up to scaler, say Jacobian is $J$: 
\[ J=
\begin{pmatrix}  x_1^{a-1} &  x_2^{a-1} & \cdots & x_n^{a-1}\\  
                 x_1^{2a-1} &  x_2^{2a-1} & \cdots &  x_n^{2a-1}\\
                 \vdots &  \vdots & \cdots & \vdots \\ 
                 x_1^{ma-1} &  x_2^{ma-1} & \cdots & x_n^{ma-1} \\
\end{pmatrix}.
\]
We can ignore the coefficients, since we are in the field of characteristic zero. 
We have $\text{ht}(I)=m$, since $I$ is generated by a regular sequence of length $m$.
Let $J^{\prime}= I_m(J)$, denotes the ideal generated by $m \times m$ minors 
of Jacobian $J$. The determinants of $m \times m$ minors of the Jacobian can be written as 
\[
J^{\prime}=\langle \; x_{i_1}^{j_1}x_{i_2}^{j_2} \cdots x_{i_m}^{j_m} \prod _{1 \leq a<b \leq m} 
(x_{i_a}^{j_a}-x_{i_b}^{j_b})  \;\rangle  \text{ for } 1 \leq i_1 <i_2<\cdots <i_m \leq n,
\]
where $j_1,j_2,\dots,j_m$ are some positive integers. Therefore 
\[ 
I+J^{\prime}=\langle p_{a}, p_{2a},\dots, p_{ma},\; x_{i_1}^{j_1}x_{i_2}^{j_2} \cdots 
x_{i_m}^{j_m} \prod _{1 \leq a<b \leq m} (x_{i_a}^{j_a}-x_{i_b}^{j_b}) \rangle. 
\]
Claim: $\sqrt{I+J^{\prime}}=( x_1,x_2,\dots,x_n).$ \\
Suppose not, that is, there exists $w \in \mathbb{P}^{n-1}$ with $w \in Z(I+J^{\prime})$.
Then the vector $w$ can have at the most $m-1$ distinct nonzero coordinates.
If $w$ has $m$ or more than $m$ distinct nonzero coordinates, then $w \notin Z(J^{\prime})$. 
Say $w$ has $v$ distinct nonzero coordinates. We can write 
\[
w=(w_1,\dots w_1, w_2,\dots,w_2, \dots, w_v,\dots,w_v,0,0,\dots,0),
\]
where $w_i$ appears $\beta_i$ times and $v\leq {m-1}$. Also $w$ should satisfy $p_{ia}$ for 
$i=1,\dots,m$. Thus, we have
\[
\beta_1 w_1^{ia} +\beta_2 w_2^{ia}+ \cdots + \beta_v w_v^{ia}=0 \text{ for $i=1,2,\dots,m.$ }
\]
This is a system of equation, which can be represented in the matrix form with $m$ 
rows and $v$ columns as
\[
 \begin{pmatrix}  1 &  1 & \cdots & 1\\  
                 w_1^a & w_2^a & \cdots & w_v^a\\
                 \vdots &  \vdots & \cdots & \vdots \\ 
                 w_1^{(m-1)a} &  w_2^{(m-1)a} & \cdots & w_v^{(m-1)a} 
\end{pmatrix}
\begin{pmatrix}  \beta_1 w_1^{a}\\  
                 \beta_2 w_2^{a}\\
                 \vdots  \\ 
                 \beta_v w_v^{a}
\end{pmatrix}
=
\begin{pmatrix}  0 \\  
                 0 \\
                 \vdots  \\ 
                 0
\end{pmatrix}.
\]
We know that neither $\beta_i=0$ nor $w_i=0$ for $i=1,\dots,v$. So $\beta_iw_i^{a}\neq0$
for $i=1,\dots,v$. We can choose the matrix say $M$ with first $v$ rows out of 
$m$ rows and look for the 
solution. The matrix $M$ is of full rank since $w_i \neq w_j$ for $i\neq j$,
so the only possible solution has to be the trivial solution. 
Hence the claim is proved. By similar argument as used in 
Theorem $\ref{main-thm-prime-ideal-any-n}$, 
we conclude that $R$ is a normal domain and $I$ is a prime ideal in $S$.
\end{proof}

For $n \geq 4$, we know that the ideal $\langle p_{1}, p_{2},\dots, p_{m} \rangle$ is prime 
in $S=\mathbb{C}[x_1,x_2,\dots,x_n]$ for all $m < n-1$, see \cite[Theorem $4.3$]{KM}. 
We will see in the following theorem that similar result holds 
for the complete symmetric polynomials and the elementary symmetric 
polynomials. 

\begin{prop}\label{n-2 h-consecutive}
Let $p_a,h_a$, and $e_a$ in the polynomial ring $S=\mathbb{C}[x_1,x_2,\dots,x_n]$, with $n \geq 4$.
Let $m < {n-1}$. Then one has
\[
\langle p_{1}, p_{2},\dots, p_{m} \rangle=\langle h_{1}, h_{2},\dots, h_{m} \rangle
=\langle e_{1}, e_{2},\dots, e_{m}\rangle.
\]
Therefore the ideals $\langle h_{1}, h_{2},\dots, h_{m} \rangle$ and 
$\langle e_{1}, e_{2},\dots, e_{m}\rangle$ are also prime in $S$.
\end{prop}
\begin{proof} It follows from simple observation in the ring of symmetric polynomials 
that the algebra generated by the power sum polynomials $p_{1}, p_{2},\dots, p_{m}$ is same as the 
algebra generated by $h_{1}, h_{2},\dots, h_{m}$, and also by $e_{1}, e_{2},\dots, e_{m}$. 
Thus one has 
\begin{eqnarray}\label{eqn11}
 \langle p_{1}, p_{2},\dots, p_{m} \rangle=\langle h_{1}, h_{2},\dots, h_{m} \rangle=
 \langle e_{1}, e_{2},\dots, e_{m}\rangle.
\end{eqnarray}
One may also conclude $(\ref{eqn11})$ from Newton's formula $(\ref{Newton's formula-2})$ 
and $(\ref{Newton's formula-1})$. By \cite[Theorem $4.3$]{KM}, $\langle p_{1}, p_{2},\dots, p_{m} \rangle$ 
is a prime ideal in $S$. Thus, we can conclude that 
$\langle h_{1}, h_{2},\dots, h_{m} \rangle$ and $\langle e_{1}, e_{2},\dots, e_{m} \rangle$ 
are also prime ideals in $S$. 
\end{proof}

\begin{rem}\label{h-1-2-5 exception}
Let $n=4$ in Proposition \ref{n-2 h-consecutive}. Then $h_1,h_2,h_a$ form a 
regular sequence in $S$ provided $h_a \notin \langle h_1,h_2 \rangle $.
By Newton's formula $(\ref{Newton's formula-2})$, 
we observe that $h_5\equiv 0 \mod (h_1,h_2)$. Hence $h_5 \in \langle h_1,h_2 \rangle $. 
We see that $h_1,h_2$ generate a prime ideal in $S$, but $h_1,h_2,h_5$ do not form a 
regular sequence in $S$. Thus, we need the condition $h_a \notin \langle h_1,h_2 \rangle $.
\end{rem} 

Computer calculations in CoCoA \cite{CoCoA} suggest that $I=\langle h_{1},h_{2m}\rangle$, 
where $m \in \mathbb{N}$, should be a prime ideal in $S=\mathbb{C}[x_1,x_2,x_3,x_4]$. 
For $m=1$, it follows from Proposition \ref{n-2 h-consecutive}. We prove for $m=2$ in 
the following example.
 
\begin{ex}\label{h-1-4} Let $S=\mathbb{C}[x_1,x_2,x_3,x_4]$ be a polynomial 
ring. Let $I=\langle h_{1},h_{4}\rangle$. Then $I$ is a prime ideal in $S$. 
\end{ex}
\begin{proof} Let $R={S}/{I}$. We compute the Jacobian of $I$, say Jacobian is $J$.
Let $J^{\prime}= I_2(J)$, denotes the ideal generated by $2 \times 2$ minors 
of Jacobian. Also $\text{ht}(I)=2$, since $I$ is generated by a regular 
sequence of length $2$. The determinants of $2 \times 2$ minors of the 
Jacobian can be written as 
\[
 J^{\prime}=\langle \; \frac{\partial h_{4}}{\partial x_i}-\frac{\partial h_{4}}{\partial x_j} 
  \;\rangle \text{ for $1 \leq i<j \leq 4$. }
\]
By Lemma \ref{h-poly} (i), we may write $J^{\prime}$ as
\[
 J^{\prime}= \langle \; (x_j-x_i)h_2+(x_j^2-x_i^2)h_1+(x_j^3-x_i^3) \rangle \text{ for $1 \leq i<j \leq 4$. }
\]
Thus, we have 
\[ 
I+J^{\prime}=\langle h_{1}, h_{4},\; (x_j-x_i)h_2+(x_j^2-x_i^2)h_1+(x_j^3-x_i^3) \rangle 
\text{ for $1 \leq i<j \leq 4$. }
\]
Claim: $\sqrt{I+J^{\prime}}=( x_1,x_2,x_3,x_4).$ \\
Suppose not, that is, there exists $w =(w_1,w_2,w_3,w_4) \in \mathbb{P}^{3}$ 
with $w \in Z(I+J^{\prime})$. We assume that none of $w_i$ is zero. Also assume that $w_i \neq w_j$ 
for $i \neq j$. Since $w$ is in $\mathbb{P}^{3}$, we can make $w_1=1$ as $w_1 \neq 0$. 
So, let $w=(1,x,y,z)$. As $w \in Z(I+J^{\prime})$, we have $h_{1}(w)=0=h_{4}(w)$, moreover 
\begin{eqnarray}\label{ext-22}
 (x_j-x_i)h_2(w)+(x_j^3-x_i^3)=0  \text{ for $1 \leq i<j \leq 4$. }
\end{eqnarray}
By $(\ref{ext-22})$, either $x_i=x_j$ or $h_2(w)=-(x_j^2+x_ix_j+x_i^2)$ for 
$1 \leq i<j \leq 4$. By assumption $x_i \neq x_j$. Thus we have $h_2(w)=-(x_j^2+x_ix_j+x_i^2)$ 
for $1 \leq i<j \leq 4$. An easy simplification shows that it is not possible. 
We may argue similarly when $w_i = w_j$ for $i \neq j$. 
Hence there is no nontrivial solution. By similar argument 
as used in Theorem $\ref{main-thm-prime-ideal-any-n}$, we conclude that 
$R$ is a normal domain and $I$ is a prime ideal.
\end{proof}

Simply note that by \cite[Lemma $2.2$ and Proposition $2.9$]{CKW}, the sequence of polynomials 
$p_{a}, p_{2a},\dots, p_{na}$ and $h_{a}, h_{2a},\dots,h_{na}$ form a regular sequence in 
$S=\mathbb{C}[x_1,x_2,\dots,x_n]$ respectively. We partially extend the above conclusion in the 
following proposition:

\begin{prop}\label{pa-ha-12n}
Let $S=\mathbb{C}[x_1,x_2,\dots,x_n]$ be a polynomial ring. Let $n,b,k \in \mathbb{N}$. 
Then the following holds:
\begin{itemize}
 \item [(i)] $p_{a},\; p_{2a},\dots,p_{(n-1)a},\;p_{b}$ form a regular sequence in $S$ if and only if $b=nak$.
\item [(ii)] $h_{a},\; h_{2a},\dots,h_{(n-1)a},\;h_{b}$ form a regular sequences in $S$ if and only if $b=nak$. 
\end{itemize}
\end{prop}

\begin{proof}
By Newton's formula $(\ref{Newton's formula-1})$, one has 
\begin{eqnarray*}\label{p-12n}
p_c= \begin{cases}
    (-1)^{k}ne_n^k \mod (p_1,p_2,\dots,p_{n-1}),  &\text{ if  $ c=nk $;}\\
    0 \mod (p_1,p_2,\dots,p_{n-1}),  &\text{ otherwise.}
  \end{cases}
\end{eqnarray*}
Clearly $p_{1}, p_{2},\dots,p_{n-1},p_{c}$ form a regular sequence in $S$ if and only if $c=nk$. 
Thus the claim (i) follows from \cite[Lemma $2.2$ ]{CKW}. By Newton's formula $(\ref{Newton's formula-2})$, one has 
 \begin{eqnarray*}\label{h-12n}
 h_c= \begin{cases}
    (-1)^{k}e_n^k \mod (h_1,h_2,\dots,h_{n-1}),  &\text{ if  $ c=nk $;}\\
    0 \mod (h_1,h_2,\dots,h_{n-1}),  &\text{ otherwise.}
    \end{cases}
\end{eqnarray*}
Clearly $h_{1}, h_{2},\dots,h_{n-1},h_{c}$ form a regular sequence in $S$ if and only if $c=nk$. 
Thus the claim (ii) also follows from \cite[Lemma $2.2$ ]{CKW}.
\end{proof}

\section{Final Remarks}
\medskip

Following question is a special case of Question \ref{question-2}:
 \begin{quest}\label{prime-ideal-p_a-p_b}
 Let $S=\mathbb{C}[x_1,x_2,\dots,x_n]$ be a polynomial ring with $n \geq 4$. 
 Let $I=\langle p_{a},p_{b} \rangle$, where $a,b \in \mathbb{N}$. For which pairs of 
 integers $a,b$, $I$ is a prime ideal in $S$.
 \end{quest}
We answer the previous question to some extent in the Theorem \ref{main-thm-prime-ideal-any-n}. 
We observe that the Theorem \ref{main-thm-prime-ideal-any-n} is still in weaker form, 
 due to arithmetic condition $q_1 > \max \{n, a\}$. In fact one can answer all the prime 
 ideals which arises by using Serre criterion for normality and vanishing sums of roots of unity, 
 by answering the following problem:

\begin{prob}\label{problem}
Let $m \in \mathbb{N}$. Consider $m$-th roots of unity in the 
field of complex number $\mathbb{C}$. For which natural numbers $n$ and $k \geq 2$, do there exist $m$-th roots 
of unity $\alpha_1,\dots,\alpha_n \in \mathbb{C}$ such that 
$\alpha_1^{k}+\alpha_2^{k}+\dots+\alpha_n^{k} \neq 0$.
\end{prob}

\medskip

Similar to Question \ref{prime-ideal-p_a-p_b}, one may also ask the following question:

\begin{quest}\label{prime-ideal-h_a-h_b}
 Let $S=\mathbb{C}[x_1,x_2,\dots,x_n]$ be a polynomial ring with $n \geq 4$. 
 Let $I=\langle h_{a},h_{b} \rangle$, where $a,b \in \mathbb{N}$. For which pairs of 
 integers $a,b$, $I$ is a prime ideal in $S$.
\end{quest}

For $n=4$ in the previous question, the computational calculations in 
CoCoA \cite{CoCoA} suggest that $I=\langle h_{1},h_{2m}\rangle$, 
where $m \in \mathbb{N}$, should be a prime ideal in $S=\mathbb{C}[x_1,x_2,x_3,x_4]$.

Recall the following definition:
\begin{defn}
Let $R=\bigoplus_{i=0}^{c}R_i,\;R_c \neq 0$ be a graded Artinian algebra. 
We say that $R$ has the strong Lefschetz property (SLP) if there exists an element 
$L\in R_1$ such that the multiplication map
 \[
  \times L^d: R_i \longrightarrow R_{i+1}
 \]
has full rank for all $0 \leq i \leq c-1$ and $1 \leq d \leq c-i$. 
We call an $L \in R_1$ with this property a strong Lefschetz element. 
\end{defn}

Let $\text{J}=\langle p_{a}, p_{a+1},\dots, p_{a+n-1} \rangle$ in the polynomial 
ring $S=K[x_1,x_2,\dots,x_n]$ over a field $K$. Then the 
Artinian ring $R=S/J$ has the SLP, see \cite[Proposition $7.1$]{Ha-Wa}. 
In the following example, we will show that 
for a complete intersection ideal $I=\langle h_{a}, h_{a+1},\dots, h_{a+n-1} \rangle$, 
the Artinian ring $R=S/I$ have the SLP.  

\begin{ex}\label{SLP-for-h-sym-ply-prop}
Let $K$ be a field of characteristic zero. Let $S=K[x_1,x_2,\dots,x_n]$ be a polynomial 
ring over $K$ with standard grading, i.e. $\deg{x_i}=1$ for all $i$. Let 
$I=\langle h_{a}, h_{a+1},\dots, h_{a+n-1} \rangle$. Then $R=S/I$ has the SLP.
\end{ex}
\begin{proof}
Consider the initial ideal of $I$ with respect to lexicographic term order, one has
\[ 
{\text{in}}(I)=\langle x_1^{a},x_2^{a+1},\dots,x_n^{a+n-1} \rangle.
\]
Stanley \cite{RS} proved that every monomial complete intersection 
  \[
   K[x_1,x_2,\dots,x_n]/{\langle x_1^{a},x_2^{a+1},\dots,x_n^{a+n-1} \rangle}
  \] 
has the SLP with $x_1+x_2+\cdots+x_n$ as a strong Lefschetz element using the 
 fact that it is isomorphic to the cohomology ring of a direct product of projective spaces 
 over the complex number field. We see that $S/in(I)$ has the SLP. Thus, by \cite[Proposition $2.9$]{AW} 
we conclude that $R$ has the SLP.
\end{proof}

\begin{quest}\label{initial-ideal} Let $S=\mathbb{C}[x_1,x_2,\dots,x_n]$ be a polynomial ring. 
Let $k \in \mathbb{N}$. Is it true that for the ideal 
$I=\langle h_{a}, h_{2a},\dots, h_{(n-1)a,h_{nak}} \rangle$, one has the initial ideal 
 \[
  {\text{in}}(I)=\langle x_1^{a},x_2^{2a},\dots,x_n^{nak} \rangle.
 \]
\end{quest}
If the answer to the previous question is positive. Then again one can construct more examples of 
Artinian ring having the SLP. In a joint work with Martino \cite{KM}, we explicitly derive several 
examples of complete intersection ideal generated by complete symmetric polynomials. Again, one can ask 
similar question for those complete intersection ideals.
\medskip

We conclude the section with one remark from the recent paper of Fr\"oberg and Shapiro \cite{Froberg-Shapiro}, 
where the authors established a connection between regular sequences of complete symmetric polynomials 
and the codimention of the Vandermonde variety.
To an arbitrary pair $(k;I)$, where $k \geq 2$ is a positive integer and $I=\{i_0<i_1< \cdots < i_{m-1} \}$, 
$m \geq k$ is a sequence of integers, Fr\"oberg and Shapiro discuss the Vandermonde variety $Vd_{k;I}^{\mathit{A}}$ 
in \cite{Froberg-Shapiro}. Under the assumption $i_0=0$ and $ \gcd (i_1,\dots,i_{m-1})=1$, 
the authors asked \cite[Problem 2]{Froberg-Shapiro}, for which pairs $(k;I)$, the 
variety $Vd_{k;I}^{\mathit{A}}$ has the expected codimention. In the first non-trivial case $k=3, m=5$, 
the authors conclude that the variety $Vd_{3;I}^{\mathfrak{A}}$ has the expected codimention 
(equal to $3$) if and only if three complete symmetric polynomials 
$h_{i_2-2},h_{i_3-2},h_{i_4-2}$ form a regular sequence in $\mathbb{C}[x_1,x_2,x_3]$. The problem of when 
three complete symmetric polynomials $h_a,h_b,h_c$, form a regular sequence in $\mathbb{C}[x_1,x_2,x_3]$ was considered in 
\cite[Conjecture $2.17$]{CKW}. In \cite[Conjecture $13$]{Froberg-Shapiro}, in a special case, 
it is mentioned that if $(a,b,c)=(1,4,3k+2), k \geq 1$, then $h_a,h_b,h_c$ neither is a regular sequence, 
nor $h_c \in (h_a,h_b)$. This will be clear from the following proposition:  

\begin{prop}\label{h-14n}
Let $S=\mathbb{C}[x_1,x_2,x_3]$ be a polynomial ring. Then $h_1,h_4,h_{n}$ form a regular sequence 
in $S$ if and only if $n=3k, k \geq 1$.
\end{prop}
\begin{proof}
By Newton's formula $(\ref{Newton's formula-2})$, one has
\begin{eqnarray*}
 h_n= \begin{cases}
    e_3^k \mod (h_1,h_4),  &\text{ if  $ n=3k $;}\\
    0 \mod (h_1,h_4),  &\text{ if  $ n=3k+1 $;}\\
    -(k+1)e_2e_3^{k}\mod (h_1,h_4),  &\text{ if $n=3k+2.$}
    \end{cases}
\end{eqnarray*}
Thus the claim follows from \cite[Theorem $2.2$]{KM}. 
\end{proof}

\end{document}